• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Moffit

    Log entry time 15:01:48 on February 8,2004

    Entry number 117051

    keyword=Reconnected Reported Helicity to Right Spectrometer

    Finished my pedestal run with the HAPPEX DAQ without Reported Helicity to the Right Spectrometer. Reconnected it (at the Counting House patch panel).

    BTW: "Pairsynch" had also been disconnected since Friday afternoon... but the Regular Spectrometer DAQ doesn't use it anyway. It's also been replaced with G0's second helicity that is sent to a second Pockels Cell in the injector. And this signal is delayed by 1, relative to the realtime helicity sent to the Pockels cell IN the laser path on the injector laser table.